Rolandos living room is 20ft.by 15ft.his ceilings are 10ft high. If a can of paint covers 150sq.ft approximately how many gallons will it take to paint the walls and ceilings
Answer by rfer(16322)

70*10=700 sqft
700+(20*15)=1000 sqft
1000/150=6.7
7 cans, can,t buy part can